# envname-P envname#1

### v1

@@ 5 Experiments
Figure 3: Better SD Potential Estimation = Better Curvature and FID. On ImgN32, convergence of dual potential $\mathbf{g}$ vs. SD-FM ($\varepsilon=0$) curvature and FID; I-FM is shown as lines. Note that curvatures of different solvers are computed on different trajectories, hence they are not comparable.
@@
Figure 4: FID vs. time needed to form a pair when training I-FM, OT-FM (varying batch sizes $n$) and SD-FM. We use $\varepsilon=0$ (SD only), $0.01,0.1$. Couplings are computed using full $d$ or PCA space. Red lines show the per-sample time $\Theta$ needed to compute the gradient of the loss for one pair. SD-FM yields significant improvements for negligible overhead.
@@
5.2 Continuous-conditional generation: super-resolution
@@ I-FM SD-FM I-FM SD-FM
As an example application to continuous-valued conditions, we apply SD-FM to train flow models for conditional sampling in image super-resolution (SR) problems. We use the CelebA dataset (Liu et al., 2015) rescaled to 64x64 resolution, containing 162k train images. We downsize them to 16x16 (4x SR) or 8x8 (8x SR) and apply Gaussian noise with standard deviation $\sigma\in\{0.1,0.2\}$ to form a condition $\mathbf{z}$ and run SD-FM with continuous data and conditions as described in Section 4.
We compare to I-FM trained by sampling $\mathbf{x}_{0}\sim\mathcal{N}(0,\operatorname{\mathbf{I}}_{d})\otimes\tilde{\mu}_{1,\mathbf{Z}}$ and from the augmented data distribution of images $\mathbf{x}$ paired with their corrupted version $\mathbf{z}$. We sample from the trained model, conditioned on corrupted images from the CelebA validation set to obtain a high-resolution reconstruction. Since the ground-truth image $\mathbf{x}$ is known, we rely on the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM, Wang et al. (2004)) to measure reconstruction accuracy. Results are summarized in Table 3 and we show sampled reconstructions from the validation set in Figure 16.
5.3 Guidance
Figure 5: Precision and recall on ImgN-32 vs. # of guidance samples $r$. $\gamma=2$, NFE =4=4, $\varepsilon=0$.
@@ 5.3 Guidance
5.4 One-Step Generation with Mean-Flow
Recently, consistency models (models that incorporate trajectory consistency constraints in their training objective) have become a popular choice for few-step generation with diffusions or flows (Song et al., 2023; Song and Dhariwal, 2024). We demonstrate that the benefits of SD-OT couplings go beyond FM by testing them on the MeanFlow (MF) model of Geng et al. (2025), noting that different consistency formulations can be unified and treated similarly (Boffi et al., 2025; Sabour et al., 2025). Figure˜6 compares SD-OT and independent couplings for training MF; we use the same setup as Geng et al. (2025) to train an unconditional DiT-XL/2 model for 300 epochs on ImgN-256x256 in the latent space.
@@ 5.4 One-Step Generation with Mean-Flow
Some aspects of the SD-FM pipeline presented here leave room for improvement. Figure 2 shows that computing the SD-OT potential to a reasonable degree of precision impacts the performance of SD-FM.
Figure 6: FID on ImgN-256 vs. NFE for mean-flows.
That SD-OT preprocessing cost occurs only once prior to FM training; it only depends on the noise distribution and data, and is typically much cheaper than the FM training cost itself. However, if SD-FM were to be applied to datasets of billions $N$ of points, this could become a challenge (though still less so than FM training). Because SD-OT is a concave maximization problem, one could leverage recipes such as batching, momentum or $\varepsilon$-tempering. For $\varepsilon>0$, categorical sampling over a $N$-sized softmax vector would be costly, which is why we highlighted the $\varepsilon=0$ case that can leverage MIPS and fast retrieval tools (both to train $\mathbf{g}$ and to pair fresh noise). Finally, having in mind dataloader efficiency, one may want to flip the tables in our sampling approach: Start from the $j$-th data point in memory, and match it with a Gaussian noise restricted (Wu and Gardner, 2024) to be in its Laguerre cell $\mathcal{L}_{j}=\{\mathbf{x}:\mathbf{x}^{T}(\mathbf{x}^{(j)}_{1}-\mathbf{x}_{1}^{(k)})+g_{j}-g_{k}\geq 0,\forall k\neq i\}$. While we have shown successful applications of the SD coupling preprocessing approach to extended FM methods such as mean-flow or guidance, SD couplings could be of course used within even more complex approaches building upon FM, such as Reflow. Our focus was on ablating the impact of the SD coupling over the independent coupling in the most widely used FM setup. Assessing how SD couplings interact with more advanced or orthogonal methods is left for future work.
@@ A.1 Properties of Semidiscrete Couplings
Plugging this back into equation A.1, we obtain
@@
$w_{t}(\mathbf{x})\coloneqq\gamma(\gamma-1)\int_{s=0}^{t}\big({\bm{v}}_{1,s}(\psi^{(\gamma)}_{s}(\mathbf{x}))-{\bm{v}}_{2,t}(\psi^{(\gamma)}_{s}(\mathbf{x}))\big)^{T}\big(\nabla\log\rho_{1,s}(\psi^{(\gamma)}_{s}(\mathbf{x}))-\nabla\log\rho_{2,s}(\psi^{(\gamma)}_{s}(\mathbf{x}))\big)\mathrm{d}s.$
@@
as $r\to\infty$. Therefore, $\rho^{(r,\gamma)}_{t}$ converges almost surely, weakly to $\rho^{(\gamma)}_{t}$. This is the sense in which the convergence of Proposition˜4 holds. We keep the argument informal since we do not study the exact concentration properties required for equation A.2 to hold.
@@
We proceed by differentiating equation A.5 against time to obtain equation A.3. We have
@@
where the last equality holds by the change of variables formula. This is exactly the weak sense of equation A.3. ∎
As a result, to prove Proposition˜5, we only need to show that $\rho^{(\gamma)}_{t}$ satisfies a PDE of the form equation A.3.
@@ A.2 Properties of the Semidual Loss
Thus, we turn to showing equation A.6. For $\varepsilon>0$,
@@
Note that at $f=f_{\mathbf{g},\varepsilon}$, $s_{f,{\mathbf{g}},{\mathbf{g}}}$ coincides with $s_{\varepsilon,{\mathbf{g}}}$ of equation 3.2. The semidual is defined as $F_{\varepsilon}({\mathbf{g}})\coloneqq\max_{f}{\mathcal{D}}_{\varepsilon}(f,{\mathbf{g}})$, and given by
@@
for $\varepsilon>0$. Thus, we derived the semidual equation $\mathrm{S}_{\varepsilon}$ for $\varepsilon>0$. Importantly, one can verify
@@
We once again derived the semidual equation $\mathrm{S}_{\varepsilon}$. Using the definition of $\pi_{0,{\mathbf{g}}}$, we have
@@
i.e. we can extend equation A.7 to $\varepsilon=0$. By Lemma˜7, for any $\varepsilon\geq 0$ we have
@@ A.3 Convergence of SGD
By the convexity of $F$, for any $\mathbf{g}$} we have $F^{*}\geq F({\mathbf{g}})+\nabla F({\mathbf{g}})^{T}({\mathbf{g}}^{*}-{\mathbf{g}})$. With this property, for any $l$ we can write
${\mathbb{E}}[\|{\mathbf{g}}_{l+1}-{\mathbf{g}}^{*}\|^{2}\,|\,{\mathbf{g}}_{t}]$ $=\|{\mathbf{g}}_{t}-{\mathbf{g}}^{*}\|^{2}-2\eta_{l}{\mathbb{E}}[\nabla f({\mathbf{g}}_{l})\,|\,{\mathbf{g}}_{l}]^{T}({\mathbf{g}}_{l}-{\mathbf{g}}^{*})+\eta_{t}^{2}{\mathbb{E}}[\|\nabla f({\mathbf{g}}_{l})\|^{2}\,|\,{\mathbf{g}}_{l}]$
$=\|{\mathbf{g}}_{l}-{\mathbf{g}}^{*}\|^{2}-2\eta_{l}\nabla F({\mathbf{g}}_{l})^{T}({\mathbf{g}}_{l}-{\mathbf{g}}^{*})+\eta_{l}^{2}{\mathbb{E}}[\|\nabla f({\mathbf{g}}_{l})\|^{2}\,|\,{\mathbf{g}}_{t}]$
@@
$F({\mathbf{g}}_{t+1})\leq F({\mathbf{g}}_{t})+\nabla F({\mathbf{g}}_{t})^{T}({\mathbf{g}}_{t+1}-{\mathbf{g}}_{t})+\frac{L}{2}\left\lVert{\mathbf{g}}_{t+1}-{\mathbf{g}}_{t}\right\rVert^{2}.$
@@
Consider the semidual $F_{\varepsilon}$ defined in equation $\mathrm{S}_{\varepsilon}$. For any cost $c:{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ and any $\varepsilon>0$, we have $\left\lVert\nabla^{2}F_{\varepsilon}\right\rVert_{\mathrm{op}}\leq 1/\varepsilon$. Furthermore, if $c(\mathbf{x},\mathbf{y})=-\mathbf{x}^{T}\mathbf{y}$ and $\mu$ admits a density, then for $\varepsilon=0$ we have $\left\lVert\nabla^{2}F_{0}\right\rVert_{\mathrm{op}}\leq 2C^{\mathrm{max}}_{\mu}/\delta$, where $\delta=\min_{i\neq j}\left\lVert\mathbf{x}^{(i)}_{1}-\mathbf{x}^{(j)}_{1}\right\rVert$, and $C^{\mathrm{max}}_{\mu}$ is the maximum $\mu$-surface area of any convex set in ${\mathbb{R}}^{d}$.
@@
$s_{0,{\mathbf{g}}}(\mathbf{x}_{0})_{i}=\mathds{1}[g_{i}+\mathbf{x}^{T}\mathbf{y}_{i}\geq g_{j}+\mathbf{x}^{T}\mathbf{y}_{j},\,\forall j],$
@@
$\nabla^{2}F_{\varepsilon}({\mathbf{g}})=\frac{1}{\varepsilon}\int[\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x})\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x})^{T}-\mathrm{diag}(\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x}))]\mathrm{d}\mu(\mathbf{x}).$
@@
where $A({\mathbf{g}},i)=\{\mathbf{x}:g_{i}+\mathbf{x}^{T}\mathbf{y}_{i}\geq g_{j}+\mathbf{x}^{T}\mathbf{y}_{j},\forall j\}$ and $\mathrm{d}S$ denotes the d1d-1-dimensional surface measure in a Euclidean space. Note that $\nabla^{2}F_{0}$ is the Laplacian of a weighted graph with weights $W_{ij}\geq 0$, therefore $\left\lVert\nabla^{2}F_{0}({\mathbf{g}})\right\rVert\leq\max_{i=1}^{N}\sum_{j\neq i}W_{ij}$. Further,
@@
As a result, the above bound also holds for ${\mathbb{E}}[\left\lVert\mathbf{g}_{t}\right\rVert]$. Combining this with equation A.8 the proof. ∎
@@ Appendix B Algorithms
We begin by detailing the optimization algorithm for solving SD-OT. If $\eta_{k}$ only depends on $k$, Algorithm˜1 is SGD on the semidual loss equation $\mathrm{S}_{\varepsilon}$. For SGD, the typical choice of learning rate is $\eta_{k}\propto 1/\sqrt{k}$ after an optional constant learning rate phase. By allowing $\eta_{k}$ to depend on the gradients, Algorithm˜1 becomes adaptive stochastic optimization on the semidual loss. We found AdaGrad (Duchi et al., 2011) to be a particularly effective choice of learning rate schedule, and provide details on the schedule of SGD and Adagrad in Appendix˜C.

### v2

@@ 5 Experiments
Figure 3: Better SD Potential Estimation = Better Curvature and FID: On ImgN32, convergence of dual potential $\mathbf{g}$ vs. SD-FM ($\varepsilon=0$) curvature and FID; I-FM is shown as lines. Note that curvatures of different solvers are computed on different trajectories, hence they are not comparable.
@@
Figure 4: FID vs. time needed to form a pair when training I-FM, OT-FM (varying batch sizes $n$) and SD-FM. We use $\varepsilon=0$ (SD only), $0.01,0.1$. Couplings are computed using full $d$ or PCA space. Red-dashed lines show the per-sample time $\Theta$ needed to compute the gradient of the loss for one pair. SD-FM yields significant improvements for a negligible overhead.
@@
5.2 Continuous-conditional generation: CelebA super-resolution
@@ I-FM SD-FM I-FM SD-FM
As an example application to continuous-valued conditions, we apply SD-FM to train flow models for conditional sampling in image super-resolution (SR) problems. We use the CelebA dataset (Liu et al., 2015) rescaled to 64x64 resolution, containing 162k train images. We downsize them to 16x16 (4x SR) or 8x8 (8x SR) and apply Gaussian noise with standard deviation $\sigma\in\{0.1,0.2\}$ to form a condition $\mathbf{z}$ and run SD-FM with continuous data and conditions as described in Section 4. We compare to I-FM trained by sampling $\mathbf{x}_{0}\sim\mathcal{N}(0,\operatorname{\mathbf{I}}_{d})\otimes\tilde{\mu}_{1,\mathbf{Z}}$ and from the augmented data distribution of images $\mathbf{x}$ paired with their corrupted version $\mathbf{z}$. We sample from the trained model, conditioned on corrupted images from the CelebA validation set to obtain a high-resolution reconstruction.
5.3 Guidance: ImgN32
Figure 5: Precision & recall on ImgN-32 vs. # of guidance samples $r$. $\gamma=2$, NFE =4=4, $\varepsilon=0$.
Since the ground-truth image $\mathbf{x}$ is known, we rely on the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM, Wang et al. (2004)) to measure reconstruction accuracy. Results are summarized in Table 3 and we show sampled reconstructions from the validation set in Figure 16.
@@ 5.3 Guidance: ImgN32
5.4 One-Step Generation with Mean-Flow: Latent Space ImgN256
Figure 6: FID on ImgN-256 vs. NFE for mean-flows.
Recently, consistency models (models that incorporate trajectory consistency regularizers in their training objective) have become a popular choice for few-step generation with diffusions or flows (Song et al., 2023; Song and Dhariwal, 2024). We demonstrate that the benefits of SD-OT couplings go beyond FM by testing them on the MeanFlow (MF) model of Geng et al. (2025), noting that different consistency formulations can be unified and treated similarly (Boffi et al., 2025; Sabour et al., 2025). Figure˜6 compares SD-OT and independent couplings for training MF; we use the same setup as Geng et al. (2025) to train an unconditional DiT-XL/2 model for 300 epochs on ImgN-256x256 in latent space and M/2 for 40 epochs. Note that SDOT in latent space may be easier than pixel space owing to the Gaussianity promoted in auto-encoder latents, see Fig.18.
@@ 5.4 One-Step Generation with Mean-Flow: Latent Space ImgN256
Some aspects of the SD-FM pipeline presented here leave room for improvement. Figure 2 shows that computing the SD-OT potential to a reasonable degree of precision impacts the performance of SD-FM. That SD-OT preprocessing cost occurs only once prior to FM training; it only depends on the noise distribution and data, and is typically much cheaper than the FM training cost itself. However, if SD-FM were to be applied to datasets of billions $N$ of points, this could become a challenge (though still less so than FM training). Because SD-OT is a concave maximization problem, one could leverage recipes such as batching, momentum or $\varepsilon$-tempering. For $\varepsilon>0$, categorical sampling over a $N$-sized softmax vector would be costly, which is why we highlighted the $\varepsilon=0$ case that can leverage MIPS and fast retrieval tools (both to train $\mathbf{g}$ and to pair fresh noise). For these reasons, and due to only very minor differences in performace, we would recommend that users use SD-FM with $\varepsilon=0$. Finally, having in mind dataloader efficiency, one may want to flip the tables in our sampling approach: Start from the $j$-th data point in memory, and match it with a Gaussian noise restricted (Wu and Gardner, 2024) to be in its Laguerre cell $\mathcal{L}_{j}=\{\mathbf{x}:\mathbf{x}^{T}(\mathbf{x}^{(j)}_{1}-\mathbf{x}_{1}^{(k)})+g_{j}-g_{k}\geq 0,\forall k\neq i\}$. While we have shown successful applications of the SD coupling preprocessing approach to extended FM methods such as mean-flow or guidance, SD couplings could be of course used within even more complex approaches building upon FM, such as Reflow. Our focus was on ablating the impact of the SD coupling over the independent coupling in the most widely used FM setup. Assessing how SD couplings interact with more advanced or orthogonal methods is left for future work.
@@ A.1 Properties of Semidiscrete Couplings
Plugging this back into (A.1), we obtain
@@
$w_{t}(\mathbf{x})\coloneqq\gamma(\gamma-1)\int_{s=0}^{t}\big({\bm{v}}_{1,s}(\psi^{(\gamma)}_{s}(\mathbf{x}))-{\bm{v}}_{2,t}(\psi^{(\gamma)}_{s}(\mathbf{x}))\big)^{\top}\big(\nabla\log\rho_{1,s}(\psi^{(\gamma)}_{s}(\mathbf{x}))-\nabla\log\rho_{2,s}(\psi^{(\gamma)}_{s}(\mathbf{x}))\big)\mathrm{d}s.$
@@
as $r\to\infty$. Therefore, $\rho^{(r,\gamma)}_{t}$ converges almost surely, weakly to $\rho^{(\gamma)}_{t}$. This is the sense in which the convergence of Proposition˜4 holds. We keep the argument informal since we do not study the exact concentration properties required for (A.2) to hold.
@@
We proceed by differentiating (A.5) against time to obtain (A.3). We have
@@
where the last equality holds by the change of variables formula. This is exactly the weak sense of (A.3). ∎
As a result, to prove Proposition˜5, we only need to show that $\rho^{(\gamma)}_{t}$ satisfies a PDE of the form (A.3).
@@ A.2 Properties of the Semidual Loss
Thus, we turn to showing (A.6). For $\varepsilon>0$,
@@
Note that at $f=f_{\mathbf{g},\varepsilon}$, $s_{f,{\mathbf{g}},{\mathbf{g}}}$ coincides with $s_{\varepsilon,{\mathbf{g}}}$ of (3.2). The semidual is defined as $F_{\varepsilon}({\mathbf{g}})\coloneqq\max_{f}{\mathcal{D}}_{\varepsilon}(f,{\mathbf{g}})$, and given by
@@
for $\varepsilon>0$. Thus, we derived the semidual ($\mathrm{S}_{\varepsilon}$) for $\varepsilon>0$. Importantly, one can verify
@@
We once again derived the semidual ($\mathrm{S}_{\varepsilon}$). Using the definition of $\pi_{0,{\mathbf{g}}}$, we have
@@
i.e. we can extend (A.7) to $\varepsilon=0$. By Lemma˜7, for any $\varepsilon\geq 0$ we have
@@ A.3 Convergence of SGD
By the convexity of $F$, for any $\mathbf{g}$} we have $F^{*}\geq F({\mathbf{g}})+\nabla F({\mathbf{g}})^{\top}({\mathbf{g}}^{*}-{\mathbf{g}})$. With this property, for any $l$ we can write
${\mathbb{E}}[\|{\mathbf{g}}_{l+1}-{\mathbf{g}}^{*}\|^{2}\,|\,{\mathbf{g}}_{t}]$ $=\|{\mathbf{g}}_{t}-{\mathbf{g}}^{*}\|^{2}-2\eta_{l}{\mathbb{E}}[\nabla f({\mathbf{g}}_{l})\,|\,{\mathbf{g}}_{l}]^{\top}({\mathbf{g}}_{l}-{\mathbf{g}}^{*})+\eta_{t}^{2}{\mathbb{E}}[\|\nabla f({\mathbf{g}}_{l})\|^{2}\,|\,{\mathbf{g}}_{l}]$
$=\|{\mathbf{g}}_{l}-{\mathbf{g}}^{*}\|^{2}-2\eta_{l}\nabla F({\mathbf{g}}_{l})^{\top}({\mathbf{g}}_{l}-{\mathbf{g}}^{*})+\eta_{l}^{2}{\mathbb{E}}[\|\nabla f({\mathbf{g}}_{l})\|^{2}\,|\,{\mathbf{g}}_{t}]$
@@
$F({\mathbf{g}}_{t+1})\leq F({\mathbf{g}}_{t})+\nabla F({\mathbf{g}}_{t})^{\top}({\mathbf{g}}_{t+1}-{\mathbf{g}}_{t})+\frac{L}{2}\left\lVert{\mathbf{g}}_{t+1}-{\mathbf{g}}_{t}\right\rVert^{2}.$
@@
Consider the semidual $F_{\varepsilon}$ defined in ($\mathrm{S}_{\varepsilon}$). For any cost $c:{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ and any $\varepsilon>0$, we have $\left\lVert\nabla^{2}F_{\varepsilon}\right\rVert_{\mathrm{op}}\leq 1/\varepsilon$. Furthermore, if $c(\mathbf{x},\mathbf{y})=-\mathbf{x}^{\top}\mathbf{y}$ and $\mu$ admits a density, then for $\varepsilon=0$ we have $\left\lVert\nabla^{2}F_{0}\right\rVert_{\mathrm{op}}\leq 2C^{\mathrm{max}}_{\mu}/\delta$, where $\delta=\min_{i\neq j}\left\lVert\mathbf{x}^{(i)}_{1}-\mathbf{x}^{(j)}_{1}\right\rVert$, and $C^{\mathrm{max}}_{\mu}$ is the maximum $\mu$-surface area of any convex set in ${\mathbb{R}}^{d}$.
@@
$s_{0,{\mathbf{g}}}(\mathbf{x}_{0})_{i}=\mathds{1}[g_{i}+\mathbf{x}^{\top}\mathbf{y}_{i}\geq g_{j}+\mathbf{x}^{\top}\mathbf{y}_{j},\,\forall j],$
@@
$\nabla^{2}F_{\varepsilon}({\mathbf{g}})=\frac{1}{\varepsilon}\int[\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x})\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x})^{\top}-\mathrm{diag}(\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x}))]\mathrm{d}\mu(\mathbf{x}).$
@@
where $A({\mathbf{g}},i)=\{\mathbf{x}:g_{i}+\mathbf{x}^{\top}\mathbf{y}_{i}\geq g_{j}+\mathbf{x}^{\top}\mathbf{y}_{j},\forall j\}$ and $\mathrm{d}S$ denotes the d1d-1-dimensional surface measure in a Euclidean space. Note that $\nabla^{2}F_{0}$ is the Laplacian of a weighted graph with weights $W_{ij}\geq 0$, therefore $\left\lVert\nabla^{2}F_{0}({\mathbf{g}})\right\rVert\leq\max_{i=1}^{N}\sum_{j\neq i}W_{ij}$. Further,
@@
As a result, the above bound also holds for ${\mathbb{E}}[\left\lVert\mathbf{g}_{t}\right\rVert]$. Combining this with (A.8) the proof. ∎
@@ Appendix B Algorithms
We begin by detailing the optimization algorithm for solving SD-OT. If $\eta_{k}$ only depends on $k$, Algorithm˜1 is SGD on the semidual loss ($\mathrm{S}_{\varepsilon}$). For SGD, the typical choice of learning rate is $\eta_{k}\propto 1/\sqrt{k}$ after an optional constant learning rate phase. By allowing $\eta_{k}$ to depend on the gradients, Algorithm˜1 becomes adaptive stochastic optimization on the semidual loss. We found AdaGrad (Duchi et al., 2011) to be a particularly effective choice of learning rate schedule, and provide details on the schedule of SGD and Adagrad in Appendix˜C.